# The Nontrivial Vacuum Structure of an Extended $t\overline{t}$ BEH (Higgs) Bound state

### v3

@@ I Introduction
A theory of a BEH boson composed of $\overline{t}t$, known as “top condensation,” was proposed in the 1990’s Yama ; BHL ; Topcolor ; NSD , deploying the Nambu-Jona-Lasinio (NJL) model NJL with renormalization group (RG) improvements BHL . The NJL model, however, is pointlike, lacking $\phi(r^{\mu})$, which leads to difficulties when there is a large hierarchy between the composite scale, $M_{0}$, and the electroweak scale $|\mu|$ (the symmetric phase mass of the BEH boson, $|\mu|=88$ GeV). The inclusion of the internal wave-function yields significant improvement in the predictions of the low energy parameters. In the limit $|\mu|<\!\!<M_{0}$ there is significant wave-function spreading of $\phi(\vec{r})$ and the resulting dilution effects dominate the low energy effective theory. This brings its predictions into concordance with experiment, virtually eliminating fine-tuning, and predicting the new mass scale of the binding interaction, $M_{0}\sim 6$ TeV, ref.(chris1 ,chris2 ) summarized in Appendix B.
@@
The issue then becomes how to avoid Lorentz non-invariance with an arbitrary input $\omega^{\mu}$, in particular, “what defines $\omega^{\mu}$?” For simple two-body bound states, where $H(X^{\mu})\sim\exp(iP_{\mu}X^{\mu})$, then $\omega^{\mu}$ can be identified with the normalized 4-momentum, $\omega^{\mu}=P^{\mu}/\sqrt{P^{2}}$ where $P^{2}=P_{\mu}P^{\mu}=\mu^{2}>0$. Hence, a two-body spherically symmetrical bound state becomes:
@@
However, there remains the question when $\mu^{2}<0$ and spontaneous symmetry breaking occurs: “How can we have a Lorentz invariant vacuum state with $P^{\mu}=0$, but nonzero $\omega^{\mu}$?” That is, “what determines $\omega^{\mu}$ in the vacuum?” Any constraint that locks $P^{\mu}$ to $\omega^{\mu}$ in the two-body case ceases to exist in the vacuum and $\omega^{\mu}$ would seem to become arbitrary. While one possibility is that $\phi(r^{\mu})\rightarrow$(constant) in the vacuum, this does not lead to a consistent solution to the SKG equation.
It would seem nonsensical to assert that the vacuum is defined by a condensate that occurred in a particular Lorentz frame with a random $\omega^{\mu}$. If, for example, $\omega^{\mu}$ is somehow associated with the local cosmic rest frame we would obtain induced Lorentz violating effects in the electroweak physics, e.g., vacuum Cerenkov radiation for all particles that receive mass from the BEH boson, top-quarks to electrons and neutrinos ColemanGlashow . We have estimated these effects in chris1 and they are suppressed as $\sim|\mu|^{2}/M_{0}^{2}$. While the limit on vacuum Cerenkov radiation for the electron is satisfied, these effects may be problematic, e.g., potentially large radiatively induced Lorentz non-invariant corrections to electrodynamics may arise at loop level Kostelecky . We therefore require a starting point in which the vacuum is manifestly Lorentz invariant, hence it must contain no preferred $\omega^{\mu}$. Yet, we evidently require nonzero $\omega^{\mu}$ to define the solutions, $\phi_{\omega}(r^{\mu})$, to implement the relative time invariance.
Hence, we propose the following solution to the vacuum problem: the vacuum is a Lorentz invariant sum over all frames of the individual solutions $\phi_{\omega}(r^{\mu})$ in each frame. We introduce a Lorentz invariant integral over $\omega^{\mu}$, leading to a novel internal wave-function for the vacuum state, $\Phi(r^{\mu})$, with (unrenormalized) $H^{\prime}(X)$.
@@
With this definition of the vacuum, $\Phi(r^{\mu})$ becomes a collective state, similar to a condensed matter system such as the BCS superconductor.
@@
Our present calculation is formally similar to the construction of the coherent condensate of Cooper pairs in a BCS superconductor Cooper BCS , differing essentially by our required Lorentz invariance. Cooper pairs are two-body bound states, (approximate) bilocal plane waves in momentum space: $\phi_{i}(\vec{r})\approx\phi_{i}\exp(\vec{k}_{i}\cdot(\vec{x}-\vec{y}))$, where the $\vec{k}_{i}$ lie in a small range immediately above the Fermi surface of the material, hence the “pairing” is an electron of momentum $\vec{k}_{i}$ with the antipodal electron $\vec{k}_{i}$. The Hamiltonian kinetic term for the pairs is then $\small\sim(1/m)\sum_{i}|\phi_{k_{i}}|^{2}k^{2}_{i}\propto N$, while the two-body attractive scattering potential (weak phonon exchange interaction) is $\small\sim-\epsilon m\sum_{i}\sum_{j}\phi^{\dagger}_{k_{i}}\phi_{k_{j}}\propto N^{2}$. Hence the weak phonon interaction is significantly enhanced by large $N\epsilon\sim 1$ and can compete with the kinetic terms to form a stable collective state which forms the condensate. The semiclassical wave-function of the ground state is then $\Phi\sim\sum_{i}^{N}\phi_{i}(\vec{r})$ where the sum spans the Fermi surface. The BCS theory is essentially a bilocal field theory. [^2] The BCS ground state is a coherent state of Cooper pairs, and specified as a product of Dirac kets that are mixtures of vacuum and pairs, $|BCS\rangle\sim\prod^{N_{F}}_{i}(u_{i}|0\rangle+v_{i}|\!\uparrow\!\vec{k}_{i},\!\downarrow\!-\vec{k}_{i}\rangle)$, where the product extends over the $N_{F}$ momenta, $k_{i}$, slightly above the Fermi surface (bounded above by the Debye frequency $\omega_{D}$). The superconductor’s properties are determined by the resulting ratio $v_{i}/u_{i}$, obtained by minimization of the Hamiltonian with, e.g., temperature effects or current dependence, etc. (See e.g., Feynstat Urbana , and related discussion of mean-field approximations GL BEC ). The condensate quantum field is bilocal, with creation operators $a^{\dagger}_{\vec{k}_{i}}$ for electrons of momenta $\vec{k}_{i}$, $\widehat{\Phi}(\vec{r})\sim\sum^{N}_{i}a^{\dagger}_{\vec{k}_{i}}a^{\dagger}_{-{\vec{k}_{i}}}\phi_{i}(\vec{r})$. Note this creates, and sums over, Cooper pairs of antipodal momenta $(\vec{k}_{i},-\vec{k}_{i})$ that span the Fermi surface. The analogue of our semiclassical bilocal field is then $\Phi(\vec{r})=\langle BCS|\widehat{\Phi}(\vec{r})|BCS\rangle\sim\sum^{N}_{i}u_{i}v_{i}^{*}\phi_{i}(\vec{r})$. The ground state (vacuum) expectation value of $\langle\Phi\rangle$ is determined by a gap equation (similar of the original formulation of the NJL model NJL ). A gap equation is equivalent to the minimization of the BEH (Higgs) potential (the “quartic interaction” of the potential is implicit in the gap equation loop, and for us the quartic coupling also arises at loop level).
@@
The vacuum emerges from the underlying theory with $H^{\prime}(X)\Phi(r^{\mu})=H(X^{\mu})\widetilde{\Phi}(r)$ where $\widetilde{\Phi}(r)=(1/\sqrt{N})\Phi(r^{\mu})$ is the classical average over all $\phi_{\omega}(r^{\mu})$ and $\langle H\rangle=v_{weak}$. The BEH boson observed at the LHC, $h(X)$ and the Nambu–Goldstone phases that become longitudinal $W^{\pm}$ and $Z^{0}$, then emerge as “excitons” of the collective state. In the broken phase we then have:
$H(X^{\mu})\widetilde{\Phi}(r^{\mu})\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}+\frac{1}{\sqrt{2}}h(X)\\ 0\end{array}\biggr)\widetilde{\Phi}(r^{\mu}),$ (7)
@@ II Implementing the Relative Time Symmetry
We begin with the notion that the internal wave-function of two-body bound state, $\phi(r^{\mu})$, must not depend upon “relative time,” i.e., is independent of $r^{0}$ in the barycentric frame, or correspondingly $r^{\prime 0}$ in any frame. This can be seen in the free field limit by kinematics if we consider a pair of equal-mass particles of 4-momenta $p_{1}$ and $p_{2}$, $p_{1}^{2}=p_{2}^{2}=m^{2}$, and a bilocal wave-function, $H(x,y)\sim\exp(ip_{1}x+ip_{2}y)$. We pass to the total momentum $P=(p_{1}+p_{2})$ and relative momentum $Q=(p_{1}-p_{2})$, and the plane waves become $\exp(iPX+iQr)$, where we define “barycentric coordinates,” $x^{\mu}=X^{\mu}+r^{\mu}$, $y^{\mu}=X^{\mu}-r^{\mu}$. Note that $P_{\mu}Q^{\mu}=p_{1}^{2}-p_{2}^{2}=0$. Therefore, in the center-of-mass frame, in which $P=(P^{0},\vec{0})$ and $Q=(0,\vec{q})$, we see that $Q^{0}=0$. This implies there is no dependence in the bilocal state on $r^{0}$ through $\exp(iQ_{0}r^{0})=1$, and likewise no dependence upon a boosted relative time $r^{\prime 0}$ in any other frame. If the particles are constituents of a bound state, then the “relative time” must decouple from the dynamics. Indeed, it is useful to think of the bound state as free field pair of particles for which the bilocal field is properly normalized, and the interaction is subsequently adiabatically switched on.
Given an arbitrary Lorentz invariant function, $\phi(r^{\mu})$, in any frame there will generally be dependence upon a relative time, $\sim\tau$. This is analogous to the gauge dependent components of a vector potential, and it is an artifact of using the bilocal field description. [^3] I am grateful to Bill Bardeen for some discussions that inspired this perspective. As described in the Introduction, the relative time, $\tau$, can be written in any given frame as $r^{\mu}=\omega^{\mu}\tau$, and implicitly requires the timelike, 4-vector, $\omega^{\mu}$, the “arrow of relative time.” Hence, eliminating dependence upon $\tau$, requires a Lorentz invariant constraint, such as $\omega_{\mu}\partial^{\mu}\phi(r)=0$. In the symmetric phase of the standard model (SM), (or for any typical two-body bound state) the BEH boson contains such a vector, i.e., the 4-momentum $P^{\mu}$ carried by $H(X^{\mu})\sim\exp(iP_{\mu}X^{\mu})$. We can therefore bootstrap $\omega^{\mu}$ to $P^{\mu}$ through a constraint relation $\omega^{\mu}\propto P^{\mu}$. For example, we can do this semiclassically by introducing Lagrange multipliers into the action, such as:
@@
We will presently focus upon the fields, $\phi_{\omega}(r^{\mu})$, as defined in eq.(2) which are invariant under the gauge transformation $r^{\mu}\rightarrow\omega^{\mu}\tau$ and thus have no dependence upon $\tau$, though an implicit dependence upon $\omega^{\mu}$ remains. Indeed, $\phi_{\omega}(r^{\mu})$ is then the analogue of a “Stueckelberg” field, e.g., a gauge field such as $B_{\mu}=A_{\mu}-\partial_{\mu}\chi$ which is invariant under $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\tau$ and $\chi\rightarrow\chi+\tau$. Note that eq.(2) satisfies the constraint equation $0=\omega^{\mu}\partial_{\mu}\phi(r^{\mu})$. In the following, $\phi(r^{\mu})$ will refer to an arbitrary Lorentz invariant scalar, while $\phi_{\omega}(r^{\mu})$ is of the (Stueckelberg) form of eq.2) with the relative time projected out.
Consider the action $S_{\phi}$ for the internal field $\phi(r^{\mu})$ that arises in natural top condensation chris1 , of eq.(87). We replace $\phi(r^{\mu})\rightarrow\phi_{\omega}(r^{\mu})$ in $S_{\phi}$,
@@
where $\partial_{\mu}=\partial/\partial r^{\mu}$ and $D_{F}(2r)$ is defined in eq.(89).
@@
Note the presence of the overall line integral, $\int\omega_{\mu}dr^{\mu}$. This line integral remains in the equation of motion since $\phi_{\omega}(r^{\mu})$ has no dependence upon $r_{\mu}\propto\omega_{\mu}$, hence the variation is constrained, $\delta\phi_{\omega}\sim\delta^{3}(r^{\mu}_{\perp})$ where $\omega_{\mu}r^{\mu}_{\perp}=0$ and does not produce a longitudinal variation, $\delta(\omega_{\mu}r^{\mu}$).
We define $Z$ by the line integral normalization chris1 ,
@@
$Z$ is defined to canonically normalize the free field pair of particles before the interaction is turned on. It is important to realize that eq.(11) is not a conventional Klein-Gordon equation due to the line integral constraint.
@@
We study its properties and solutions in ref.chris1 .
In the solution to the SKG equations with the eigenvalue, $\mu^{2}$, we see upon integrating by parts that $S_{\phi}\rightarrow\mu^{2}$, and the full bound state action then becomes (in the pointlike limit of the interaction):
@@
$\mu^{2}$ is the physical mass of the bound state. The Yukawa coupling, $g_{Y}$ and quartic coupling $\lambda$ are defined in eqs.(85-90) in Appendix B. We emphasize that, while we have evaluated the action in the particular $\omega^{\mu}$ frame in eq.(13), this is just a calculation in a simplifying frame in the overall Lorentz invariant action eq.(11). The result eq.(17) holds in any frame.
@@
The solution of the SKG equation for $\phi(r)$ can be obtained approximately analytically, or by numerical integration chris1 ; chris2 . At short distances $\phi(r)\sim\phi(0)$, and extends at large distances in the rest frame to, $\phi(r)\sim ce^{-|\mu|r}/r$, where $|\mu|<\!\!M_{0}$ and we are near critical coupling. The solution is normalized as in eq.(14), which dilutes the value of $\phi(0)\sim\sqrt{|\mu|/M_{0}}$ and suppresses the Yukawa coupling $g_{Y}\propto\phi(0)$ and $\lambda\propto g_{Y}^{4}\propto|\phi(0)|^{4}$.

### v4

@@ I Introduction
A theory of a BEH boson composed of $\overline{t}t$, known as “top condensation,” was proposed in the 1990’s Yama ; BHL ; Topcolor ; NSD , deploying the Nambu-Jona-Lasinio (NJL) model NJL BHL . The NJL model, however, is pointlike, lacking $\phi(r^{\mu})$, which leads to difficulties when there is a large hierarchy between the composite scale, $M_{0}$, and the electroweak scale $|\mu|$ (the symmetric phase mass of the BEH boson, $|\mu|=88$ GeV). The inclusion of the internal wave-function yields significant improvement in the predictions of the low energy parameters. In the limit $|\mu|<\!\!<M_{0}$ there is significant wave-function spreading of $\phi(\vec{r})$ and the resulting dilution effects dominate the low energy effective theory. This brings its predictions, including the quartic coupling $\lambda$, into concordance with experiment, virtually eliminating fine-tuning and predicting the new mass scale of the binding interaction, $M_{0}\sim 6$ TeV, ref.(chris1 ,chris2 ) summarized in Appendix B.
@@
The issue then becomes “what defines $\omega^{\mu}$?” For simple two-body bound states, where $H(X^{\mu})\sim\exp(iP_{\mu}X^{\mu})$, then $\omega^{\mu}$ can be identified with the normalized 4-momentum, $\omega^{\mu}=P^{\mu}/\sqrt{P^{2}}$, where the bound state mass is $P^{2}=P_{\mu}P^{\mu}=\mu^{2}>0$. Hence, a two-body spherically symmetrical bound state becomes:
@@
However, there remains the question when $\mu^{2}<0$ and spontaneous symmetry breaking occurs: “How can we have a Lorentz invariant vacuum state with $P^{\mu}=0$, but nonzero $\omega^{\mu}$?” That is, “what determines $\omega^{\mu}$ in the vacuum?” Any constraint that locks $P^{\mu}$ to $\omega^{\mu}$ ceases to exist in the vacuum and $\omega^{\mu}$ would seem to become arbitrary. One possibility is that $\phi(r^{\mu})\rightarrow$(constant) in the vacuum, but this does not lead to a consistent solution to the SKG equation with a potential that depends upon $r^{\mu}$.
It would seem nonsensical to assert that the vacuum is defined by a condensate that occurred in a particular Lorentz frame with a random $\omega^{\mu}$. If, for example, $\omega^{\mu}$ is somehow associated with the local cosmic rest frame we would obtain induced Lorentz violating effects in the electroweak physics, e.g., vacuum Cerenkov radiation for all particles that receive mass from the BEH boson, top-quarks to electrons and neutrinos ColemanGlashow . We have estimated these effects in chris1 and, for a particle of mass $m_{f}$, they are suppressed as $\sim m_{f}|\mu|^{2}/M_{0}^{2}$. While the limit on vacuum Cerenkov radiation for the electron is satisfied, these effects are likely problematic, e.g., potentially large radiatively induced Lorentz non-invariant corrections to electrodynamics may arise at loop level Kostelecky . We therefore require a starting point in which the vacuum is manifestly Lorentz invariant, hence it must contain no preferred $\omega^{\mu}$. Yet, we evidently require nonzero $\omega^{\mu}$ to define the solutions, $\phi_{\omega}(r^{\mu})$, to implement the relative time invariance.
Hence, we arrive at the following solution to the vacuum problem: the vacuum is a Lorentz invariant sum over all frames of the individual solutions $\phi_{\omega}(r^{\mu})$ in each frame. We introduce a Lorentz invariant integral over $\omega^{\mu}$, leading to a novel internal wave-function for the vacuum state, $\Phi(r^{\mu})$, with (unrenormalized) $H^{\prime}(X)$.
@@
With this definition of the vacuum, $\Phi(r^{\mu})$ becomes a collective state, similar to a condensed matter system such as the BCS superconductor. Note the return of relative time, $r^{0}$, but we will see that $r^{\mu}$ can now be integrated out of the final action, yielding Lorentz invariant higher dimension operators suppressed as $\sim M_{0}^{-2p}$ for $p\geq 1$.
@@
Our present calculation is formally similar to the construction of the coherent condensate of Cooper pairs in a BCS superconductor Cooper BCS , differing essentially by our required Lorentz invariance. Cooper pairs are two-body bound states, (approximate) bilocal plane waves in momentum space: $\phi_{i}(\vec{r})\approx\phi_{i}\exp(\vec{k}_{i}\cdot(\vec{x}-\vec{y}))$, where the $\vec{k}_{i}$ lie in a small range immediately above the Fermi surface of the material, hence the “pairing” is an electron of momentum $\vec{k}_{i}$ with the antipodal electron $\vec{k}_{i}$. The Hamiltonian kinetic term for the pairs is then $\small\sim(1/m)\sum_{i}|\phi_{k_{i}}|^{2}k^{2}_{i}\propto N$, while the two-body attractive scattering potential (weak phonon exchange interaction) is $\small\sim-\epsilon m\sum_{i}\sum_{j}\phi^{\dagger}_{k_{i}}\phi_{k_{j}}\propto N^{2}$. Hence the weak phonon interaction is significantly enhanced by large $N\epsilon\sim 1$ and can compete with the kinetic terms to form a stable collective state which forms the condensate. The semiclassical wave-function of the ground state is then $\Phi\sim\sum_{i}^{N}\phi_{i}(\vec{r})$ where the sum spans the Fermi surface. The BCS theory is essentially a bilocal field theory. [^2] The BCS ground state is a coherent state of Cooper pairs that is a product of Dirac kets that are mixtures of the Fermi-degenerate groundstate and pairs, $|BCS\rangle\sim\prod^{N_{F}}_{i}(u_{i}|0\rangle+v_{i}|\!\uparrow\!\vec{k}_{i},\!\downarrow\!-\vec{k}_{i}\rangle)$, where the product extends over the $N_{F}$ momenta, $k_{i}$, slightly above the Fermi surface (bounded above by the Debye frequency $\omega_{D}$). The superconductor’s properties are determined by the resulting ratio $v_{i}/u_{i}$, obtained by minimization of the Hamiltonian with, e.g., temperature effects or current dependence, etc. (See e.g., Feynstat Urbana , and related discussion of mean-field approximations GL BEC ). The condensate quantum field is bilocal, with creation operators $a^{\dagger}_{\vec{k}_{i}}$ for electrons of momenta $\vec{k}_{i}$, $\widehat{\Phi}(\vec{r})\sim\sum^{N}_{i}a^{\dagger}_{\vec{k}_{i}}a^{\dagger}_{-{\vec{k}_{i}}}\phi_{i}(\vec{r})$. Note this creates, and sums over, Cooper pairs of antipodal momenta $(\vec{k}_{i},-\vec{k}_{i})$ that span the Fermi surface. The analogue of our semiclassical bilocal field is then $\Phi(\vec{r})=\langle BCS|\widehat{\Phi}(\vec{r})|BCS\rangle\sim\sum^{N}_{i}u_{i}v_{i}^{*}\phi_{i}(\vec{r})$. The ground state (vacuum) expectation value of $\langle\Phi\rangle$ is determined by a gap equation (similar of the original formulation of the NJL model NJL ). A gap equation is equivalent to the minimization of the BEH (Higgs) potential (the “quartic interaction” of the potential is implicit in the gap equation loop, and for us the quartic coupling also arises at loop level).
@@
The vacuum emerges from the underlying theory as $\langle H\rangle{\Phi}(r)$ where $\Phi(r^{\mu})$ is the coherent sum over all $\phi_{\omega}(r^{\mu})$ and $\langle H\rangle=v_{weak}$. The BEH boson observed at the LHC, $h(X)$ and the Nambu–Goldstone phases that become longitudinal $W^{\pm}$ and $Z^{0}$, then emerge as “excitons” of the collective state. In the broken phase we then have:
$H(X^{\mu}){\Phi}(r^{\mu})\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}+\frac{1}{\sqrt{2}}h(X)\\ 0\end{array}\biggr){\Phi}(r^{\mu}),$ (7)
@@ II Implementing the Relative Time Symmetry
We begin with the notion that the internal wave-function of two-body bound state, $\phi(r^{\mu})$, must not depend upon “relative time,” i.e., is independent of $r^{0}$ in the barycentric frame, or correspondingly $r^{\prime 0}$ in any frame. This can be seen in the free field limit by kinematics if we consider a pair of equal-mass particles of 4-momenta $p_{1}$ and $p_{2}$, $p_{1}^{2}=p_{2}^{2}=m^{2}$, and a bilocal wave-function, $H(x,y)\sim\exp(ip_{1}x+ip_{2}y)$. We pass to the total momentum $P=(p_{1}+p_{2})$ and relative momentum $Q=(p_{1}-p_{2})$, and the plane waves become $\exp(iPX+iQr)$, where we define “barycentric coordinates,” $x^{\mu}=X^{\mu}+r^{\mu}$, $y^{\mu}=X^{\mu}-r^{\mu}$. Note that $P_{\mu}Q^{\mu}=p_{1}^{2}-p_{2}^{2}=0$. Therefore, in the center-of-mass frame, in which $P=(P^{0},\vec{0})$ and $Q=(0,\vec{q})$, we see that $Q^{0}=0$. This implies there is no dependence in the bilocal state on $r^{0}$ through $\exp(iQ_{0}r^{0})=1$, and likewise no dependence upon a boosted relative time $r^{\prime 0}$ in any other frame. If the particles are constituents of a bound state, then the “relative time” must decouple from the dynamics. Indeed, it is useful to think of the bound state as free field pair of particles for which the bilocal field is properly normalized, and the interaction is subsequently adiabatically switched on, maintaining this kinematic constraint. The free particle states then “flow” to become the bound states.
Given an arbitrary Lorentz invariant function, $\phi(r^{\mu})$, in any frame there will generally be dependence upon a relative time, $\sim\tau$. This is analogous to the gauge dependent components of a vector potential, and it is an artifact of using the bilocal field description. [^3] I am grateful to Bill Bardeen for some discussions that inspired this perspective. As described in the Introduction, the relative time, $\tau$, can be written in any given frame as $r^{\mu}=\omega^{\mu}\tau$, and implicitly requires the timelike, 4-vector, $\omega^{\mu}$, the “arrow of relative time.” Hence, eliminating dependence upon $\tau$, requires a Lorentz invariant constraint, such as $\omega_{\mu}\partial^{\mu}\phi(r)=0$.
In the symmetric phase of the standard model (SM), (or for any typical two-body bound state) the BEH boson contains such a vector, i.e., the 4-momentum $P^{\mu}$ carried by $H(X^{\mu})\sim\exp(iP_{\mu}X^{\mu})$. We can therefore bootstrap $\omega^{\mu}$ to $P^{\mu}$ through a constraint relation $\omega^{\mu}\propto P^{\mu}$. For example, we can do this semiclassically by introducing Lagrange multipliers into the action, such as:
@@
We will presently focus upon the fields, $\phi_{\omega}(r^{\mu})$, as defined in eq.(2) which are invariant under the gauge transformation $r^{\mu}\rightarrow\omega^{\mu}\tau$ and thus have no dependence upon $\tau$, though an implicit dependence upon $\omega^{\mu}$ remains. [^4] Indeed, $\phi_{\omega}(r^{\mu})$ is then the analogue of a “Stueckelberg” field, e.g., a gauge field such as $B_{\mu}=A_{\mu}-\partial_{\mu}\chi$ which is invariant under $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\tau$ and $\chi\rightarrow\chi+\tau$. Note that eq.(2) satisfies the constraint equation $0=\omega^{\mu}\partial_{\mu}\phi(r^{\mu})$. In the following, $\phi(r^{\mu})$ will refer to an arbitrary Lorentz invariant scalar, while $\phi_{\omega}(r^{\mu})$ is of the (Stueckelberg) form of eq.(2) with the relative time projected out.
Consider the action, $S_{\phi}$ of eq.(88), for the internal field $\phi(r^{\mu})$ that arises in natural top condensation chris1 . We replace $\phi(r^{\mu})\rightarrow\phi_{\omega}(r^{\mu})$ in $S_{\phi}$,
@@
where $\partial_{\mu}=\partial/\partial r^{\mu}$ and $D_{F}(2r)$ is the potential as defined in eq.(90).
@@
Note the presence of the overall line integral, $\int\omega_{\mu}dr^{\mu}$. This line integral remains in the equation of motion since $\phi_{\omega}(r^{\mu})$ has no dependence upon $r_{\mu}\propto\omega_{\mu}$, hence the variation is constrained, $\delta\phi_{\omega}\sim\delta^{3}(r^{\mu}_{\perp})$ where $\omega_{\mu}r^{\mu}_{\perp}=0$ and does not produce a longitudinal variation, $\delta(\omega_{\mu}r^{\mu}$). It is important to realize that eq.(11) is not a conventional Klein-Gordon equation due to the line integral constraint. Here we define $Z$ by the line integral normalization chris1 ,
@@
$Z$ is thus defined to canonically normalize the free field pair of particles before the interaction is turned on, where $\phi_{\omega}$ is a dimensionless field and normalized as:
$M_{0}^{4}Z\int\!\!d^{4}r|\phi(r^{\mu})|^{2}=1.$ (13)
@@
We have discussed its properties and solutions in ref.chris1 .
In the solution to the SKG equation with the eigenvalue, $\mu^{2}$, we see upon integrating by parts that $S_{\phi}\rightarrow\mu^{2}$, and the full bound state action, given in eq.87 of Appendix B, then becomes:
@@
$\mu^{2}$ is the physical mass of the bound state. Note that in the pointlike limit of the interaction, the internal fields $\phi_{\omega}(\vec{r})$ remain extended. The Yukawa coupling, $g_{Y}$ and quartic coupling $\lambda$ are defined in eqs.(25-91) in Appendix B. We emphasize that, while we have evaluated the action in the particular $\omega^{\mu}$ frame in eq.(14), this is just a calculation in a simplifying frame in the overall Lorentz invariant action eq.(11). The result eq.(18) holds in any frame.
@@
The solution of the SKG equation for $\phi(r)$ can be obtained approximately analytically, or by numerical integration chris1 ; chris2 . At short distances $\phi(r)\sim\phi(0)$, and extends at large distances in the rest frame to, $\phi(r)\sim ce^{-|\mu|r}/r$, where $|\mu|<\!\!M_{0}$ and we are near critical coupling. The solution is normalized as in eq.(15), which dilutes the value of $\phi(0)\sim\sqrt{|\mu|/M_{0}}$ and suppresses the Yukawa coupling $g_{Y}\propto\phi(0)$. Inputting $g_{Y}\approx 1$ yields $M_{0}\approx 6$ TeV, and we find $\lambda\propto g_{Y}^{4}$, as given by loops in eq.(91, remarkably close to the expermental value.